OpenStudy (anonymous):

We have 12 letters- 4 A's 3 B's 2 C's and D,E&F The are to be arranged in a row such that the C's are separated from one another. The total number of ways this can be done is-

OpenStudy (amistre64):

the rest of the letters might as well be all the same ....

OpenStudy (amistre64):

12 spots, 2 cs, and need to be separte

OpenStudy (amistre64):

cxcxxxxxxxxx cxxcxxxxxxxx .... there are 9 of those :) move over by 1 xcxcxxxxxxx tthere are 8 of those 9+8+7+6 .... is what im thinking

OpenStudy (amistre64):

err, starts with 10 ways eh

OpenStudy (anonymous):

_A_A_A_A_B_B_B_D_E_F_ the As Bs and D,E,F can be arranged in \[\frac{10!}{4!*3!}\] ways...(1) now we can fill 12 remaining spots with 3Cs.....is 12C3 ways.....(2) multiply (1)(2)

OpenStudy (anonymous):

sorry (2) its in 12C2 ways...

OpenStudy (amistre64):

my idea takes me too: 10(10 + 1)/2 = 5(11) = 55 ways ....

OpenStudy (anonymous):

answer in the book is 1386000 ... I'm not quite getting there

OpenStudy (amistre64):

im of course ignoring that some letters are positioned in other manners, so it prolly a bit low lol

OpenStudy (amistre64):

number of different ways to arrange the nonC letter ... times 55 ways to put in the cs afterwards ... maybe

ganeshie8 (ganeshie8):

you can also try solving like this : 1) find total number of ways of arranging the given 12 letters 2) find total number of ways of arranging 12 letters such that 2Cs are together 3) subtract 2 from 1

ganeshie8 (ganeshie8):

1) 12! / 4!3!2! 2) 11! / 4!3! 3) ?

OpenStudy (amistre64):

\[\frac{n}{a!b!c!...}\]

OpenStudy (amistre64):

\[\frac{10!}{4!3!2!}*55\] lets see if the wolf likes it lol

OpenStudy (amistre64):

693 000 aint what the book says, so ugh!!

ganeshie8 (ganeshie8):

@amistre64 logic is correct... its \(\frac{10!}{4!3!}*55\) 2! shouldnt be there in the bottom as we fixed c's to start with !!

OpenStudy (anonymous):

@amistre64 u should remove the 2!....in the denominator...